OpenStudy (anonymous):

principal branch of log(1-i)

OpenStudy (anonymous):

if someone can go step by step i would really appreciate it

OpenStudy (valpey):

pv log(z) = Log(z) = ln|z| + i(Arg(z))

OpenStudy (anonymous):

i have absolutely no clue on this problem

OpenStudy (valpey):

As in solutions to tan(x) = y, there are many values for x, but just one principal value (the one that solves Arctan(y) = x).

OpenStudy (valpey):

z=1-i |z| = sqrt(2) Arg(z) = -pi/4